# Enhancing Sample Selection Against Label Noise by Cutting Mislabeled Easy Examples

### v2

@@ 4.1 Preliminary Presentation of Effectiveness
We first provide empirical evidence to verify the effectiveness of Early Cutting. Using CIFAR-10 with 40% various synthetic label noise and ResNet-18 as the backbone, we compared our proposed Early Cutting with loss-based and dynamic-based sample selection methods. Table 1 shows that Early Cutting consistently achieved the highest test accuracy across all noise types. Although Early Cutting and the dynamic-based method selected training subsets with similar noise rates, our approach’s better performance indicates that focusing on filtering specific harmful mislabeled examples improves selection quality. The last row shows the number of additional samples filtered by Early Cutting and the high percentage of mislabeled samples among them, proving its effectiveness at identifying and removing noisy labels. As intuition, more challenging noise types result in more mislabeled samples being removed, leading to larger performance gains. Detailed settings in Appendix B.
@@ 4.2 Comparison with the Competitors
Competitors. We compare our approach with several methods: robust loss functions including GCE (Zhang and Sabuncu, 2018) and Student Loss (Zhang et al., 2024b); robust training methods, including Co-teaching (Han et al., 2018) and CSGN (Lin et al., 2024b); and sample selection methods, including Me-Momentum (Bai and Liu, 2021), Self-Filtering (Wei et al., 2022), VOG (Agarwal et al., 2022), Late Stopping (Yuan et al., 2023), Misdetect (Deng et al., 2024) and RLM (Li et al., 2024a).
Table 2: Test performance (mean±$\pm$±std) of each approach using ResNet-18 on CIFAR-10.
@@
Table 3: Test performance (mean±$\pm$±std) of each approach using ResNet-34 on CIFAR-100.
@@
Table 4: Test performance (mean±$\pm$±std) of each approach using ResNet-18 and 34 on CIFAR-N.
@@
Datasets and implementation. We conducted experiments on several benchmark datasets to evaluate our proposed method compare with above competitors. For synthetic noise experiments, we used CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009), adding symmetric and instance-dependent label noise at rates of 20% and 40% following standard protocols (Bai and Liu, 2021; Yuan et al., 2023). We split 10% noisy trianing data for validation. For real-world noisy labels, we utilized CIFAR-N (Wei et al., 2021b), as well as the large-scale WebVision dataset (Li et al., 2017). Following previous work (Lin et al., 2024b; Li et al., 2024a), we used the first 50 classes of the WebVision dataset and validated on both the WebVision validation set and the ILSVRC12 (Russakovsky et al., 2015) validation set. We further confirmed the scalability of our proposed method on the full ImageNet-1K (Deng et al., 2009) with 40% synthetic symmetric label noise. Training was performed using SGD with a momentum of 0.9 and a weight decay of $5\times 10^{-4}$. The initial learning rate was set to 0.1 and decayed using a cosine annealing schedule. Models were trained (for the final iteration) for 300 epochs on the CIFAR datasets, for 200 epochs on WebVision, and for 150 epochs on full ImageNet-1k. We re-implemented all competitor methods with consistent settings (unless otherwise specified). Detailed settings are provided in Appendix B.
@@
Figure 5: Sensitivity analysis of hyperparameters on CIFAR-10 and CIFAR-100 with 40% symmetric label noise. In each subfigure, the left plot shows test accuracy versus thresholds for the large loss, high confidence, and low gradient norm criteria, scaled by factors of $\frac{1}{4}$, $\frac{1}{2}$, $1$, $2$, and $4$. The right plot shows test accuracy versus Early Cutting rate $\gamma$ set to $n^{1/3}$, $n^{1/2}$, $n$, $n^{2}$, and $n^{3}$, where $\gamma\geq 1$.
Discussions on experimental results. As shown in Tables 2, 3, 4, and 5, our proposed Early Cutting method consistently achieves outstanding performance across various datasets and noise conditions. On standard benchmarks like CIFAR, it attains the highest test accuracy regardless of the type of label noise—symmetric, instance-dependent, or real-world—and across different noise rates. Notably, it performs exceptionally well on CIFAR-100, which has a larger number of classes, indicating strong robustness in handling label noise in fine-grained classification tasks. Without further fine-tuning of hyperparameters, Early Cutting also achieves significant performance improvements on large-scale datasets such as WebVision and full ImageNet-1k. This demonstrates the practicality and scalability of our method in handling challenging scenarios. By iteratively selecting confident samples and removing harmful mislabeled easy examples, our method helps the model learn from reliable data while avoiding overconfidence in early-learned samples. Appendix B.8 provides a training time evaluation, demonstrating that our proposed Early Cutting method, executed once per training round, incurs an additional computational overhead of less than one percent of the total training duration. This is significantly lower than the base computational cost associated with selecting the initial confident subset $\mathcal{D}^{s}$ in each training round.
4.3 Further Analysis
Semi-supervised learning. To further evaluate the effectiveness of our Early Cutting method, we integrated it with the MixMatch semi-supervised learning framework (Berthelot et al., 2019), resulting in Early Cutting+. We treat the confident samples obtained from sample selection as labeled data, and the samples removed from training in the fully supervised setting as unlabeled data. We compared Early Cutting+ with advanced SSL-based LNL methods, including CORES2∗ (Cheng et al., 2020), DivideMix (Li et al., 2020), and ELR+ (Liu et al., 2020); additionally, we compared with the latest sample selection methods integrating SSL, SFT+ (Wei et al., 2022) and RLM+ (Li et al., 2024a). Some baseline results are taken from Wei et al. (2022). As shown in Table 6, our Early Cutting+ achieves the highest test accuracy on both CIFAR-10 and CIFAR-100 under 50% symmetric and 40% instance-dependent label noise, surpassing previous methods. These results underscore the capability of Early Cutting in selecting high-quality subsets of training samples.
Sensitivity analysis. We conducted sensitivity analyses to evaluate the robustness of our Early Cutting method. As shown in Figure 5, our method achieves optimal test accuracy on both CIFAR-10 and CIFAR-100 with 40% symmetric label noise when using the same default thresholds (also used for WebVision and full ImageNet-1k) for identifying samples with large loss, high confidence, low gradient norm, and early cutting rate. Notably, even when the hyperparameters vary over a wide range, our method exhibits minimal sensitivity, with only slight ($<1\%$) performance degradation. Results indicate that our method is robust and effective across different datasets without requiring extensive hyperparameter tuning. A further ablation study is presented in Appendix B.9.
@@ 5 Conclusion
In this paper, we uncovered an oversight in existing methods for learning with noisy labels by demonstrating that not all mislabeled examples harm the model’s performance equally. We identified a specific subset termed Mislabeled Easy Examples (MEEs)—mislabeled samples that the model learns early and that significantly mislead the training process. To address this issue, we proposed Early Cutting, a counter-intuitive sample selection strategy that recalibrates the confident subset by leveraging the model’s later training state—which is typically considered unreliable—to effectively filter out MEEs. This work provides a practical solution for learning with noisy labels and advances the understanding of how different mislabeled samples affect deep learning models.
@@ Appendix A Related Work
Learning with Noisy Labels (LNL) has been an active research area in recent years (Wang et al., 2022; Li et al., 2024b; Wang et al., 2024; Xia et al., 2023, 2022; Gong et al., 2023; Luo et al., 2024; Baek et al., 2024; Englesson and Azizpour, 2024), with numerous methods proposed to mitigate the impact of label noise on deep neural networks (DNNs). Formally, let $X$ denote the input space, and let $Y=\{1,2,\dots,C\}$ be the set of clean labels. Consider the clean data distribution $P(X,Y)$, from which clean samples $(\mathbf{x},y)$ are drawn. In practice, we often have access only to a training dataset with potentially noisy labels:
@@
where $\mathbf{x}_{i}\in X$ and $\tilde{y}_{i}\in Y$ are observed noisy labels. The aim is to learn a robust classifier $f:X\rightarrow Y$ parameterized by $\theta$, which performs well on clean test data drawn from the distribution $P(X,Y)$. The noise process is typically modeled using a noise transition matrix $T\in\mathbb{R}^{C\times C}$, defined as:
$T_{ij}=P(\tilde{y}=j\mid y=i),\quad\text{for }i,j\in\{1,2,\dots,C\},$ (7)
@@
$P(\tilde{y}\mid\mathbf{x})=\sum_{k=1}^{C}T_{k\tilde{y}}\,P(y=k\mid\mathbf{x}).$ (8)
@@
Learning dynamics reaearch for sample selection. The intriguing generalization ability of modern DNNs has motivated extensive studies on their learning dynamics, which in turn has inspired a series of sample selection criteria using in Eq.(11) based on these dynamics. Studies have revealed that hard and mislabeled examples are typically learned during the later stages of training (Arpit et al., 2017; Song et al., 2019b, 2021; Maini et al., 2022; Bai et al., 2021; Lin et al., 2024a). This empirical observation has led to the development of various training-time metrics to quantify the “hardness” of examples, such as learning speed (Maini et al., 2022; Jiang et al., 2021). These metrics have inspired LNL approaches that leverage learning dynamics to select clean samples. Methods like Self-Filtering (Wei et al., 2022), FSLT & SSFT (Maini et al., 2022), SELFIE (Song et al., 2019a), and RoCL (Zhou et al., 2021) adopt criteria to identify clean samples based on their learning dynamics. The success of learning dynamics-based sample selection criteria in identifying high-confidence clean samples has driven researchers to further refine these strategies. By identifying a larger subset of clean samples for model training, the generalization performance of the trained model can be improved. (Xia et al., 2021) discovered that using loss alone to select CHEs is suboptimal. RLM (Li et al., 2024a) obtain robust loss estimation for noisy samples.
@@
Hard label noise. Various forms of hard label noise have been studied, including asymmetric noise (Scott et al., 2013), instance-dependent noise (Xia et al., 2020b), natural noise (Wei et al., 2021b), adversarially crafted labels (Zhang et al., 2024a), open-set noise (Wei et al., 2021a), and subclass-dominant noise (Bai et al., 2023). These noise types are designed from the perspective of the labels, aiming to simulate challenging real-world scenarios or malicious attacks. Recent work has also explored the impact of label noise in specific data distributions. For instance, H2E (Yi et al., 2022) and TABASCO (Lu et al., 2023) focus on the challenges posed by label noise in long-tailed distributions, where minority classes are more susceptible to mislabeling. NoiseCluster (Bai et al., 2023) introduces the concept of subclass-dominant label noise, where mislabeled examples dominate at least one subclass, leading to suboptimal classifier performance.
@@
Our contributions. In contrast to prior studies that mainly focus on different types of label noise or sample selection based on learning dynamics, our work offers a fresh perspective by re-examining sample selection methods that rely on a model’s early learning stages. We demonstrate that some samples hidden among those considered “confident” are, in fact, the most harmful when mislabeled. Specifically, we systematically investigate the detrimental impact of Mislabeled Easy Examples (MEEs)—mislabeled samples that are correctly predicted by the model early in the training process. This insight challenges the conventional assumptions of existing methods, which often prioritize samples learned early in training as being clean. Our findings highlight the need for a more cautious approach when selecting samples based on early learning confidence. By adopting a refined sample selection criterion that accounts for the potential harm of MEEs, we can seamlessly integrate this approach with existing sample selection method (Yuan et al., 2023) to further boost it performance.
Appendix B Detailed Settings
B.1 Datasets

### v3

@@ 4.1 Preliminary Presentation of Effectiveness
We first provide empirical evidence to verify the effectiveness of Early Cutting. Using CIFAR-10 with 40% various synthetic label noise and ResNet-18 as the backbone, we compared our proposed Early Cutting with loss-based and dynamic-based sample selection methods. Table 1 shows that Early Cutting consistently achieved the highest test accuracy across all noise types. Although Early Cutting and the dynamic-based method selected training subsets with similar noise rates, our approach’s better performance indicates that focusing on filtering specific harmful mislabeled examples improves selection quality. The last row shows the number of additional samples filtered by Early Cutting and the high percentage of mislabeled samples among them, proving its effectiveness at identifying and removing noisy labels. As intuition, more challenging noise types result in more mislabeled samples being removed, leading to larger performance gains. Detailed settings in Appendix D.
@@ 4.2 Comparison with the Competitors
Competitors. We compare our approach with several methods: robust loss functions including GCE (Zhang and Sabuncu, 2018) and Student Loss (Zhang et al., 2024b); robust training methods, including Co-teaching (Han et al., 2018) and CSGN (Lin et al., 2024b); and sample selection methods, including Me-Momentum (Bai and Liu, 2021), Self-Filtering (Wei et al., 2022), VOG (Agarwal et al., 2022), Late Stopping (Yuan et al., 2023), Misdetect (Deng et al., 2024) and RLM (Li et al., 2024).
[…]
@@
[…]
@@
[…]
@@
Datasets and implementation. We conducted experiments on several benchmark datasets to evaluate our proposed method compare with above competitors. For synthetic noise experiments, we used CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009), adding symmetric and instance-dependent label noise at rates of 20% and 40% following standard protocols (Bai and Liu, 2021; Yuan et al., 2023). We split 10% noisy trianing data for validation. For real-world noisy labels, we utilized CIFAR-N (Wei et al., 2021b), as well as the large-scale WebVision dataset (Li et al., 2017). Following previous work (Lin et al., 2024b; Li et al., 2024), we used the first 50 classes of the WebVision dataset and validated on both the WebVision validation set and the ILSVRC12 (Russakovsky et al., 2015) validation set. We further confirmed the scalability of our proposed method on the full ImageNet-1K (Deng et al., 2009) with 40% synthetic symmetric label noise. Training was performed using SGD (Robbins and Monro, 1951) with a momentum (Rumelhart et al., 1986) of 0.9 and a weight decay (Krogh and Hertz, 1991) of $5\times 10^{-4}$. The initial learning rate was set to 0.1 and decayed using a cosine annealing schedule. Models were trained (for the final iteration) for 300 epochs on the CIFAR datasets, for 200 epochs on WebVision, and for 150 epochs on full ImageNet-1k. We re-implemented all competitor methods with consistent settings (unless otherwise specified). For all experiments reporting mean±std, the results are the average and standard deviation of three trials using different random seeds. Detailed settings are provided in Appendix D.
Discussions on experimental results. As shown in Tables 2, 3, 4, and 5, our proposed Early Cutting method consistently achieves outstanding performance across various datasets and noise conditions. On standard benchmarks like CIFAR, it attains the highest test accuracy regardless of the type of label noise—symmetric, instance-dependent, or real-world—and across different noise rates. Notably, it performs exceptionally well on CIFAR-100, which has a larger number of classes, indicating strong robustness in handling label noise in fine-grained classification tasks. Without further fine-tuning of hyperparameters, Early Cutting also achieves significant performance improvements on large-scale datasets such as WebVision and full ImageNet-1k. This demonstrates the practicality and scalability of our method in handling challenging scenarios. By iteratively selecting confident samples and removing harmful mislabeled easy examples, our method helps the model learn from reliable data while avoiding overconfidence in early-learned samples. Furthermore, to demonstrate the transferability of our method to different model architectures, we conducted additional experiments on the transformer-based TinyViT (Vaswani et al., 2017; Wu et al., 2022), which are detailed in Appendix D.9. Appendix D.7 provides a training time evaluation, demonstrating that our proposed Early Cutting method, executed once per training round, incurs an additional computational overhead of less than one percent of the total training duration. This is significantly lower than the base computational cost associated with selecting the initial confident subset $\mathcal{D}^{s}$ in each training round.
4.3 Further Analysis
Semi-supervised learning. To further evaluate the effectiveness of our Early Cutting method, we integrated it with the MixMatch (Berthelot et al., 2019) semi-supervised learning framework, resulting in Early Cutting+. We treat the confident samples obtained from sample selection as labeled data, and the samples removed from training in the fully supervised setting as unlabeled data. The confident set (labeled data) and the non-confident set (unlabeled data) are identified once and remain fixed throughout the subsequent MixMatch training. We compared Early Cutting+ with advanced SSL-based LNL methods, including CORES2∗ (Cheng et al., 2020), DivideMix (Li et al., 2020), and ELR+ (Liu et al., 2020); additionally, we compared with the latest sample selection methods integrating SSL, SFT+ (Wei et al., 2022) and RLM+ (Li et al., 2024). Some baseline results are taken from Wei et al. (2022). As shown in Table 6, our Early Cutting+ achieves the highest test accuracy on both CIFAR-10 and CIFAR-100 under 50% symmetric and 40% instance-dependent label noise, surpassing previous methods. These results underscore the capability of Early Cutting in selecting high-quality subsets of training samples.
@@ 4.3 Further Analysis
Figure 5: Sensitivity analysis of hyperparameters on CIFAR-10 and CIFAR-100 with 40% symmetric label noise. In each subfigure, the left plot shows test accuracy versus thresholds for the large loss, high confidence, and low gradient norm criteria, scaled by factors of $\frac{1}{4}$, $\frac{1}{2}$, 11, 22, and $4$. The right plot shows test accuracy versus Early Cutting rate $\gamma$ set to $n^{1/3}$, $^{1/2}, nn$, $n^{2}$, and $n^{3}$, where $\gamma\geq 1$.
Sensitivity analysis. We conducted sensitivity analyses to evaluate the robustness of our Early Cutting method. As shown in Figure 5, our method achieves optimal test accuracy on both CIFAR-10 and CIFAR-100 with 40% symmetric label noise when using the same default thresholds (also used for WebVision and full ImageNet-1k) for identifying samples with large loss, high confidence, low gradient norm, and early cutting rate. Notably, even when the hyperparameters vary over a wide range, our method exhibits minimal sensitivity, with only slight ($1\%$) performance degradation. Results indicate that our method is robust and effective across different datasets without requiring extensive hyperparameter tuning. A further ablation study is presented in Appendix D.8.
@@ 5 Conclusion
In this paper, we uncovered an oversight in existing methods for learning with noisy labels by demonstrating that not all mislabeled examples harm the model’s performance equally. We identified a specific subset termed Mislabeled Easy Examples (MEEs)—mislabeled samples that the model learns early and that significantly mislead the training process. To address this issue, we proposed Early Cutting, a counter-intuitive sample selection strategy that recalibrates the confident subset by leveraging the model’s later training state, which is typically considered unreliable, to effectively filter out MEEs. This work provides a practical solution for learning with noisy labels and advances the understanding of how different mislabeled samples affect deep learning models.
Limitation. We acknowledge several limitations that suggest avenues for future work. First, our empirical validation is confined to visual classification tasks, and the method’s applicability to other domains such as natural language processing or tabular data remains to be explored. Second, while our work provides strong empirical evidence for the detrimental impact of MEEs, this justification is primarily observational; a formal theoretical analysis explaining why these specific examples disproportionately harm generalization would further strengthen our claims. Third, as an iterative strategy, our method introduces computational overhead by requiring re-evaluation of samples and computation of metrics like input gradients at later training stages (see Appendix D.7 for a detailed analysis). Finally, potential failure modes may arise in extreme data scenarios. In datasets with severe label noise or significant class imbalance (e.g., long-tailed distributions), the distinction between MEEs and clean hard examples from rare classes could become ambiguous. This poses a risk of inadvertently filtering out valuable samples from minority classes, which could potentially introduce fairness concerns by biasing the final model against underrepresented groups.
Acknowledgments
The authors would like to thank the anonymous reviewers for their insightful and constructive comments. The authors are also grateful to the Area Chairs for their diligent work. During the preparation of this paper, the authors utilized large language models to enhance the clarity of the writing and to generate code for training and visualization. TL is partially supported by the following Australian Research Council projects: FT220100318, DP220102121, LP220100527, LP220200949. BH was supported by RGC Young Collaborative Research Grant No. C2005-24Y and RGC General Research Fund No. 12200725.
@@ Appendix A Related Work
Learning with Noisy Labels (LNL) has been an active research area in recent years (Wang et al., 2022; Gong et al., 2023; Baek et al., 2024; Englesson and Azizpour, 2024; Huang et al., 2024), with numerous methods proposed to mitigate the impact of label noise on deep neural networks (DNNs). Formally, let $X$ denote the input space, and let $\mathcal{Y}=\{1,2,\dots,K\}$ be the set of possible labels. Let $Y$ be the random variable for the clean label and $\tilde{Y}$ be the random variable for the observed noisy label, both taking values in $\mathcal{Y}$. Consider the clean data distribution $P(X,Y)$, from which clean samples $(\mathbf{x},y)$ are drawn. In practice, we often have access only to a training dataset with potentially noisy labels:
@@
where $\mathbf{x}_{i}\in X$ and $\tilde{y}_{i}\in\mathcal{Y}$ are observed noisy labels. The aim is to learn a robust classifier $f:X\rightarrow\mathcal{Y}$ parameterized by $\theta$, which performs well on clean test data drawn from the distribution $P(X,Y)$. The noise process is typically modeled using a noise transition matrix $T\in\mathbb{R}^{K\times K}$, defined as:
$T_{ij}=P(\tilde{Y}=j\mid Y=i),\quad\text{for }i,j\in\mathcal{Y},$ (8)
@@
$P(\tilde{Y}=\tilde{y}\mid X=\mathbf{x})=\sum_{k\in\mathcal{Y}}T_{k\tilde{y}}\,P(Y=k\mid X=\mathbf{x}).$ (9)
@@
Learning dynamics reaearch for sample selection. The intriguing generalization ability of modern DNNs has motivated extensive studies on their learning dynamics, which in turn has inspired a series of sample selection criteria using in Eq.(12) based on these dynamics. Studies have revealed that hard and mislabeled examples are typically learned during the later stages of training (Arpit et al., 2017; Song et al., 2019b, 2021; Maini et al., 2022; Li et al., 2023; Lin et al., 2024a; Hong et al., 2024). This empirical observation has led to the development of various training-time metrics to quantify the “hardness” of examples, such as learning speed (Maini et al., 2022; Jiang et al., 2021) and gradient variance. For instance, Agarwal et al. (2022) proposed Variance of Gradients (VoG) to estimate sample difficulty based on the temporal variability of gradient norms, while Novak et al. (2018) analyzed generalization through input-output Jacobian norms, connecting sensitivity in input space to learning robustness. These metrics have inspired LNL approaches that leverage learning dynamics to select clean samples. Methods like Self-Filtering (Wei et al., 2022), FSLT & SSFT (Maini et al., 2022), SELFIE (Song et al., 2019a), and RoCL (Zhou et al., 2021) adopt criteria to identify clean samples based on their learning dynamics. The success of learning dynamics-based sample selection criteria in identifying high-confidence clean samples has driven researchers to further refine these strategies. By identifying a larger subset of clean samples for model training, the generalization performance of the trained model can be improved. (Xia et al., 2021) discovered that using loss alone to select CHEs is suboptimal. RLM (Li et al., 2024) obtain robust loss estimation for noisy samples.
@@
Hard label noise. Various forms of hard label noise have been studied (Lin et al., 2023), including asymmetric noise (Scott et al., 2013), instance-dependent noise (Xia et al., 2020b), natural noise (Wei et al., 2021b), adversarially crafted labels (Zhang et al., 2024a), open-set noise (Wei et al., 2021a), and subclass-dominant noise (Bai et al., 2023). These noise types are designed from the perspective of the labels, aiming to simulate challenging real-world scenarios or malicious attacks. Recent work has also explored the impact of label noise in specific data distributions. For instance, H2E (Yi et al., 2022) and TABASCO (Lu et al., 2023) focus on the challenges posed by label noise in long-tailed distributions, where minority classes are more susceptible to mislabeling. NoiseCluster (Bai et al., 2023) introduces the concept of subclass-dominant label noise, where mislabeled examples dominate at least one subclass, leading to suboptimal classifier performance.
@@
Our contributions. In contrast to prior studies that mainly focus on different types of label noise or sample selection based on learning dynamics, our work offers a fresh perspective by re-examining sample selection methods that rely on a model’s early learning stages. We demonstrate that some samples hidden among those considered “confident” are, in fact, the most harmful when mislabeled. Specifically, we systematically investigate the detrimental impact of Mislabeled Easy Examples (MEEs)—mislabeled samples that are correctly predicted by the model early in the training process. This insight challenges the conventional assumptions of existing methods, which often prioritize samples learned early in training as being clean. Our findings highlight the need for a more cautious approach when selecting samples based on early learning confidence. By adopting a refined sample selection criterion that accounts for the potential harm of MEEs, we can seamlessly integrate this approach with existing sample selection method (Yuan et al., 2023) to further boost it performance. Furthermore, our proposed method conceptually linking to Sharpness-Aware Minimization (SAM) (Foret et al., 2021), but applied in the input space to identify stable yet mislearned examples.
Appendix B Quantitative Analysis of MEE Harm with Influence Functions
To provide a more direct and quantitative measure of the harm caused by Mislabeled Easy Examples (MEEs), we conducted an analysis using influence functions (Koh and Liang, 2017). This method allows us to calculate the impact score of different training samples on the model’s performance (i.e., loss) on a clean, held-out validation set. A positive score indicates a harmful influence, while a negative score indicates a beneficial one. This analysis, performed using the pytorch_influence_functions library on CIFAR-10 with 40% instance-dependent noise, parallels the experiment in Figure 2. We compared three distinct sample categories: (1) MEEs, defined as the first 4,000 mislabeled samples learned during training; (2) Clean Easy Examples, the first 4,000 clean samples learned; and (3) Mislabeled Hard Examples, the last 4,000 mislabeled samples learned. The results provide direct, quantitative proof of our central claim. While all mislabeled examples exhibited a harmful positive influence, the average harm of MEEs was significantly greater than that of Mislabeled Hard Examples, with a mean influence score of 4.96 versus 3.01 (and a median of 4.48 vs. 3.26). This demonstrates at a microscopic level that MEEs are far more detrimental to the model. In contrast, the Clean Easy Examples showed a mean negative influence score of -2.22 (median -1.89), confirming their beneficial impact on generalization. This analysis further substantiates our claim in Section 2.2 regarding the unique and disproportionate harm caused by MEEs.
Appendix C Discussion on Broader Applicability
While the empirical validation in this paper is focused on visual classification, the core concepts of Mislabeled Easy Examples (MEEs) and Early Cutting are not inherently limited to this domain. We briefly discuss the potential translation of these ideas to other areas. In Natural Language Processing (NLP), an MEE could manifest as a text sample with strong but misleading keywords. For instance, a sarcastic positive review such as, "Wow, I can’t believe how awful the service was," being mislabeled as "Positive". A model might quickly and confidently learn this incorrect association due to the presence of words like “Wow” and “can’t believe”, which often appear in positive contexts. The Early Cutting criteria could be conceptually adapted by, for example, measuring the gradient norm with respect to the input word embeddings to gauge stability. Similarly, in regression tasks, an MEE might be a data point where a specific feature has a strong, simple, and local linear relationship with the target value, but this relationship is globally spurious or an artifact of noise. The model would likely fit this deceptive pattern early in training. Our criteria could be conceptually translated by identifying samples that the later-stage model predicts with high confidence (e.g., in a value range far from the given label, thus incurring high loss) and stable input gradients. Though these examples are conceptual, they suggest that the underlying principle of identifying and removing confidently mislearned easy examples (MEEs) could be a generalizable strategy for enhancing model robustness across diverse machine learning domains.
Appendix D Detailed Settings
D.1 Datasets